Please help! I need to identify whether these four equations are a parabola, line, circle, or ellipse. I think I have some of them worked out, but please help me.
* y^2-4x-2y-6=0 parabola?
* y^2+x^2-9=0 ellipse?
* 2x^2+3y-8x+6=0 parabola?
* 3y^2+5y^2-6y-20x-36=0 circle?
Thank you so much your a life saver.

Answer by stanbon(75887)

Please help! I need to identify whether these four equations are a parabola, line, circle, or ellipse. I think I have some of them worked out, but please help me.
* y^2-4x-2y-6=0 parabola?
one squared variable, so parabola
----------------------------------------
* y^2+x^2-9=0 ellipse?
two squared variables with equal coefficients, so circle
-------------------------------------------------------------
* 2x^2+3y-8x+6=0 parabola?
one squared variable, so parabola
------------------------------------------------------------
* 3y^2+5y^2-6y-20x-36=0 circle?
two squared variables with unequal coefficients, so ellipse
===============================================================
cheers,
Stan H.
=============